# A Survey of Artificial Intelligence in Gait-Based Neurodegenerative Disease Diagnosis

### v1

@@ 3.2.6. Distribution of Sample Sizes
Table 1. Existing studies (P-1 to P-124) of AI-assisted PD diagnosis based on gait. We summarize the input data, main AI model, experimental sample size, and reported accuracy (range) of each study. Quality scores ($S_{Q}$) that evaluate the advancement of AI models, comprehensiveness of comparison, and sufficiency of samples are computed.
@@
P-1 (tahir2012parkinson, ) 2012 PD Senosr+Vision
@@
P-2 (kugler2013automatic, ) 2013 PD Sensor EMG SVM 10 90.0 0.33
P-3 (klucken2013unbiased, ) 2013 PD Sensor IMUs AdaBoost 173 81.0 0.56
P-4 (manap2013parkinsonian, ) 2013 PD Senosr+Vision
@@
P-5 (zhang2013pathological, ) 2013 PD Sensor vGRF SVM 62 81.5-83.4 0.56
P-6 (tripoliti2013automatic, ) 2013 PD Sensor IMUs RF, DT, NB 16 96.1 0.33
P-7 (khorasani2014hmm, ) 2014 PD Sensor Force sensitive resistor LS-SVM, HMM 31 90.3 0.67
P-8 (arora2014high, ) 2014 PD Sensor IMUs RF 20 98.0 0.33
P-9 (wahid2015classification, ) 2015 PD Sensor vGRF KFD, BA, KNN, SVM, RF 49 82.0-92.6 0.44
P-10 (tupa2015motion, ) 2015 PD Vision RGB Camera NN 51 97.2 0.67
P-11 (alkhatib2015gait, ) 2015 PD Sensor vGRF KNN 166 83.0 0.56
P-12 (prochazka2015bayesian, ) 2015 PD Senosr+Vision Kinect (Camera+Depth sensor)
@@
P-13 (perumal2016gait, ) 2016 PD Sensor IMUs+Force sensitive resistor LDA, SVM, ANN 166 83.3-90.0 0.56
P-14 (zeng2016parkinson, ) 2016 PD Sensor vGRF
@@
P-15 (wang2016recognizing, ) 2016 PD Sensor Force sensitive resistor Neural networks 165 87.9 1.00
P-16 (ertuugrul2016detection, ) 2016 PD Sensor vGRF
@@
P-17 (jane2016q, ) 2016 PD Sensor Force sensitive resistor Q-BTDNN 166 90.9-92.2 1.00
P-18 (cuzzolin2017metric, ) 2017 PD Sensor IMUs KNN 580 80.8-90.2 0.56
P-19 (abujrida2017smartphone, ) 2017 PD Sensor IMUs (Smartphone) RF 50 87.0 0.44
P-20 (alam2017vertical, ) 2017 PD Sensor IMUs+vGRF SVM, KNN, RF, DT 47 93.6 0.67
P-21 (joshi2017automatic, ) 2017 PD Sensor vGRF SVM 31 90.3-100 0.67
P-22 (djuric2017selection, ) 2017 PD Sensor Walkway system (GAITRite) RF, RBF-SVM 80 85.0 0.44
P-23 (pham2017tensor, ) 2017 PD Sensor vGRF Tensor decomposition 165 100.0 0.56
P-24 (wu2017measuring, ) 2017 PD Sensor vGRF GLRA, SVM 58 82.8-84.5 0.44
P-25 (fernandes2018artificial, ) 2018 PD Sensor IMUs MLP, DBNs 45 93.5-94.5 0.56
P-26 (urcuqui2018exploring, ) 2018 PD Sensor Kinect LR, NB, RF, DT 60 61.0-82.0 0.44
P-27 (abdulhay2018gait, ) 2018 PD Sensor vGRF SVM 166 94.8 0.56
P-28 (caramia2018imu, ) 2018 PD Sensor IMUs NB, LDA, k-NN, DT, SVM 50 68.9-75.6 0.44
P-29 (mileti2018measuring, ) 2018 PD Sensor IMUs+Force sensitive resistor
@@
P-30 (rovini2018comparative, ) 2018 PD Sensor IMUs SVM, RF, NB 90 78.0-96.7 0.67
P-31 (pham2017tensor, ) 2018 PD Sensor vGRF Tensor model 165 86.2 0.78
P-32 (aich2018validation, ) 2018 PD Sensor IMUs SVM, KNN, DT, NB 51 89.1 0.67
P-33 (mittra2018classification, ) 2018 PD Sensor vGRF LR, DT, RF, SVM, KNN 166 93.1 0.56
P-34 (chen2018detecting, ) 2018 PD Sensor vGRF SVM, PSO 135 87.1-95.7 0.56
P-35 (aich2018machine, ) 2018 PD Vision Camera+Reflective markers
@@
P-36 (krajushkina2018gait, ) 2018 PD Sensor Kinect Fisher’s Score (FS) 40 75.0-85.0 0.33
P-37 (acsurouglu2018parkinson, ) 2018 PD Sensor vGRF
@@
P-38 (elden2018computer, ) 2018 PD Sensor vGRF RBF-SVM 64 90.6 0.67
P-39 (zhao2018hybrid, ) 2018 PD Sensor vGRF LSTM, CNN 166 93.9-98.6 0.89
P-40 (buongiorno2019low, ) 2019 PD Vision Camera (Kinect) SVM, ANN 30 89.4 0.44
Table 2. Continued table of Table 1 for AI-assisted PD diagnosis based on gait.
P-41 (rehman2019comparison, ) 2019 PD Sensor IMUs SVM, RF 196 0.805-0.878 (AUC) 0.56
P-42 (khoury2019data, ) 2019 PD Sensor vGRF
@@
P-43 (reyes2019lstm, ) 2019 PD Vision Camera (Kinect) Conv LSTM 182 64.0-83.0 0.67
P-44 (DBLP:conf/hicss/RastegariAA19, ) 2019 PD Sensor IMUs
@@
P-45 (guo2019method, ) 2019 PD Sensor EEG RF, SVM, KNN 41 85.7 0.44
P-46 (vasquez2018multimodal, ) 2019 PD Sensor IMUs CNN 84 97.6 0.56
P-47 (andrei2019parkinson, ) 2019 PD Sensor Force sensitive resistor SVM 166 88.9-100.0 0.78
P-48 (rehman2019selecting, ) 2019 PD Sensor Walkway system (GAITRite) RF, SVM, LR 303 97.0 0.56
P-49 (ricciardi2019using, ) 2019 PD Sensor+Vision
@@
P-50 (zeng2019classification, ) 2019 PD Sensor vGRF NN 166 98.8 0.78
P-51 (alharbi2020genetic, ) 2020 PD Sensor Force sensitive resistor
@@
P-52 (el2020deep, ) 2020 PD Sensor vGRF DNN 165 98.7 0.89
P-53 (perez2020identification, ) 2020 PD Sensor IMUs+Force sensitive resistor SVM 10 97.4-98.8 0.33
P-54 (alkhatib2020machine, ) 2020 PD Sensor vGRF Discriminant analysis 47 95.0 0.44
P-55 (abujrida2020machine, ) 2020 PD Sensor IMUs (Smart Phone) RF 456 85.5-95.0 0.56
P-56 (veeraragavan2020parkinson, ) 2020 PD Sensor vGRF NN 166 97.4 0.89
P-57 (hughes2019models, ) 2020 PD Sensor Force sensitive resistor
@@
P-58 (xia2019dual, ) 2020 PD Sensor vGRF CNN, LSTM 166 98.0 0.89
P-59 (perez2020identification, ) 2020 PD Sensor IMUs+Force sensitive resistor
@@
P-60 (som2020unsupervised, ) 2020 PD Sensor IMUs MLP, SVM 34 68.9 0.67
P-61 (rehman2020turning, ) 2020 PD Sensor IMUs PLS-DA 93 97.9 0.44
P-62 (nair2020predicting, ) 2020 PD Sensor IMUs LR 16 93.6 0.33
P-63 (moon2020classification, ) 2020 PD Sensor vGRF NN, SVM, KNN, DT, RF 567 49.0-61.0 (F1 score) 0.67
P-64 (gong2020novel, ) 2020 PD Vision Camera Mask R-CNN 301 97.3 0.67
P-65 (balaji2020supervised, ) 2020 PD Sensor vGRF SVM, DT 166 69.9-99.4 0.78
P-66 (butt2020biomechanical, ) 2020 PD Sensor IMUs Bidirectional-LSTM 114 82.4 0.67
P-67 (zhang2020deep, ) 2020 PD Sensor IMUs CNN 2804 0.86 (AUC) 0.67
P-68 (yurdakul2020detection, ) 2020 PD Sensor vGRF ANN 166 98.3 0.89
P-69 (balaji2021automatic, ) 2021 PD Sensor vGRF LSTM 64 98.6 0.89
P-70 (pham2021classification, ) 2021 PD Sensor vGRF Bi-LSTM 116 91.2-100.0 0.89
P-71 (balaji2021data, ) 2021 PD Sensor vGRF KNN, NB, SVM 165 88.5-98.8 0.78
P-72 (williamson2021detecting, ) 2021 PD Sensor IMUs GMM 380 0.69 (AUC) 0.56
P-73 (mirelman2021detecting, ) 2021 PD Sensor IMUs
@@
P-74 (ma2023explainable, ) 2021 PD Sensor Force sensitive resistor CNN, XGBoost 166 97.3-98.4 0.78
P-75 (alharthi2020gait, ) 2021 PD Sensor vGRF CNN 166 92.7-98.3 1.00
P-76 (yuan2021accelerating, ) 2021 PD Others Electronic medical record LR, RNN 28216 0.874 (AUC) 0.78
P-77 (park2021classification, ) 2021 PD Vision
@@
P-79 (shalin2021prediction, ) 2021 PD Sensor vGRF LSTM 11 84.5 0.67
P-80 (guayacan2021visualising, ) 2021 PD Vision Camera CNN 22 94.9 0.44
Table 3. Continued table of 2 for AI-assisted PD diagnosis based on gait.
P-81 (ghaderyan2021inter, ) 2021 PD Sensor vGRF
@@
P-82 (nguyen2021classification, ) 2021 PD Sensor vGRF NN 166 84.8 1.00
P-83 (zheng2021svm, ) 2021 PD Sensor IMUs SVM 300 96.7 0.56
P-84 (ahamed2021spatiotemporal, ) 2021 PD Sensor vGRF CNN 166 97.0 0.67
P-85 (salimi2023type, ) 2022 PD Sensor vGRF KNN, Fuzzy neural network 165 97.6 1.00
P-86 (carvajal2022classification, ) 2022 PD Sensor IMUs CNN, GRU 134 83.7-92.7 0.67
P-87 (mehra2022design, ) 2022 PD Sensor vGRF RF 165 94.5-98.8 0.78
P-88 (li2022detecting, ) 2022 PD Sensor vGRF LR, SVM, DT, KNN 306 81.0-85.0 0.56
P-89 (ogata2022foot, ) 2022 PD Sensor IMUs SVM 115 92.7-96.3 0.56
P-90 (rehman2022investigating, ) 2022 PD Sensor IMUs SVM, RF 99 64.1-72.7 0.44
P-91 (trabassi2022machine, ) 2022 PD Sensor IMUs SVM, DT, RF, KNN, MLP 128 74.0-86.0 0.56
P-92 (munoz2022machine, ) 2022 PD Vision Camera (Kinect) LR, SVM, DT, NB, RF 60 81.8-84.5 0.44
P-93 (yang2022pd, ) 2022 PD Sensor IMUs PD-ResNet 457 95.5 1.00
P-94 (guayacan2022quantification, ) 2022 PD Sensor Force sensitive resistor SVM, RF 22 99.6 0.33
P-95 (goni2022smartphone, ) 2022 PD Sensor IMUs (Smart phone) RF 1397 73.0 0.56
P-96 (doi:10.1080/20476965.2022.2125838, ) 2022 PD Vision Camera
@@
P-97 (acsurouglu2022deep, ) 2022 PD Sensor vGRF CNN, RF 166 99.5 1.00
P-98 (olmos2022riemannian, ) 2022 PD Vision Camera
@@
P-99 (ingelse2022personalised, ) 2022 PD Sensor IMUs CNN, NN 17 91.0-92.0 0.56
P-100 (filtjens2022automated, ) 2022 PD Sensor Retro-reflective markers
@@
P-101 (carvajal2022effective, ) 2022 PD Sensor IMUs XGBoost, SVM, RF 134 81.0-88.3 0.56
P-102 (chatzaki2022can, ) 2022 PD Sensor+Vision Camera+Force sensitive resistor AdaBoost, Extra Trees, RF 44 88.0 0.44
P-103 (ferreira2022machine, ) 2022 PD Sensor+Vision Camera+Force sensitive resistor
@@
P-104 (brand2022gait, ) 2022 PD Sensor IMUs DCNN, LSTM 30 83.3 0.78
P-105 (atri2022deep, ) 2022 PD Sensor IMUs (Smart phone)
@@
P-106 (zhao2021multimodal, ) 2022 PD Sensor+Vision Camera (Kinect)+vGRF+IMUs
@@
P-107 (mendoza2022local, ) 2022 PD Vision Camera CNN, SVM 22 96.6 0.44
P-108 (nino2023parkinsonian, ) 2022 PD Vision Camera
@@
P-109 (zeng2022slowfast, ) 2022 PD Vision Camera (Smart phone) SlowFast GCN network 68 77.1-87.5 0.89
P-110 (li2022abnormal, ) 2022 PD Sensor Force sensitive resistor+IMUs RF, LR, Gradient Boosting 22 97.3 0.33
P-111 (wang2022gait, ) 2022 PD Sensor vGRF SVM 166 98.2 0.78
P-112 (goh2022gait, ) 2022 PD Sensor vGRF SVM, ANN 48318 93.0 0.56
P-113 (kour2022vision, ) 2022 PD Sensor+Vision
@@
P-114 (beigi2022classification, ) 2022 PD Sensor IMUs MLP, KNN, RF, SVM 30 100.0 0.56
P-115 (baby2017parkinsons, ) 2022 PD Sensor vGRF ANN 166 64.4-86.8 0.67
P-116 (divyashree2022ai, ) 2023 PD Sensor vGRF 2D CNN 166 98.0 0.89
P-117 (beigi2023freezing, ) 2023 PD Sensor IMUs
@@
P-118 (uchitomi2023classification, ) 2023 PD Sensor IMUs CNN 90 82.8-86.4 0.55
P-119 (cai2023specific, ) 2023 PD Sensor IMUs PCA, SVM 280 88.0-90.0 0.55
P-120 (sun2023novel, ) 2023 PD Sensor Force sensitive resistor
@@
P-121 (zhang2023wm, ) 2023 PD Vision Camera (Smart phone)
@@
P-122 (dotov2023coordination, ) 2023 PD Sensor IMUs SVM 78 82.0 0.44
P-123 (dong2023static, ) 2023 PD Sensor IMUs
@@
P-124 (han2023automatic, ) 2023 PD Sensor IMUs Non-linear model 53 84.9 0.44
Table 4. Existing studies (C-1 to C-19) that include combinatorial diagnosis of different NDs (PD, ALS, HD) based on gait. We summarize the input data, main AI model, experimental sample size, reported accuracy (range), and quality score of each study.
@@
C-1 (daliri2012automatic, ) 2012 PD, ALS, HD Sensor Force sensitive resistor SVM 64 90.6 0.44
C-2 (dutta2013hybrid, ) 2013 PD, ALS, HD Sensor vGRF Elman’s recurrent neural network (ERNN) 63 90.6 0.78
C-3 (xia2015classification, ) 2015 PD, ALS, HD Sensor Force sensitive resistor SVM, RF, MLP, KNN 64 96.6-100.0 0.44
C-4 (pratiher2017towards, ) 2017 PD, ALS, HD Sensor vGRF SVM, EC, DT, KNN 62 79.3 0.56
C-5 (najafabadian2018neurodegenerative, ) 2017 PD, ALS, HD Sensor vGRF Adaboost 64 87.3-92.3 0.67
C-6 (athisakthi2017statistical, ) 2018 PD, ALS, HD Sensor vGRF RF, SVM 64 90.6-96.9 0.44
C-7 (klomsae2018string, ) 2018 PD, ALS, HD Sensor Force sensitive resistor fusszy KNN 64 98.4 0.44
C-8 (yan2020classification, ) 2020 PD, ALS, HD Sensor Force sensitive resistor Random forest 64 79.3-91.7 0.67
C-9 (begum2020recognition, ) 2020 PD, ALS, HD Vision Camera Discriminant Analysis, RF, Multi-SVM, KNN 64 92.2 0.44
C-10 (nam2020development, ) 2020 PD, ALS, HD Sensor Force sensitive resistor SVM, KNN 64 99.8-100.0 0.44
C-11 (lin2020evaluation, ) 2020 PD, ALS, HD Sensor vGRF CNN 64 98.4-100.0 0.56
C-12 (beyrami2020robust, ) 2020 PD, ALS, HD Sensor vGRF
@@
C-13 (yilmaz2020evaluation, ) 2020 PD, ALS, HD Sensor Force sensitive resistor KNN 64 90.5 0.44
C-14 (setiawan2021identification, ) 2021 PD, ALS, HD Sensor vGRF CNN, SVM 64 97.4-100.0 0.78
C-15 (zhao2021multimodal, ) 2022 PD, ALS, HD Sensor+Vision
@@
C-16 (setiawan2022development, ) 2022 PD, ALS, HD Sensor vGRF AlextNet CNN 63 96.4 0.78
C-17 (tobar2022petri, ) 2022 PD, ALS, HD Sensor vGRF RF, classification trees 64 69.6-95.1 0.67
C-18 (torres2023using, ) 2023 PD, ALS, HD Sensor Force sensitive resistor

### v2

@@ 4.2.3 A-DL Models for PD Diagnosis
Table 2: Existing studies (P-1 to P-124) of AI-assisted PD diagnosis based on gait. We summarize input data, main AI model(s), experimental sample size, and accuracy (range) of each study. Quality scores ($S_{Q}$) that evaluate advancement of AI models, comprehensiveness of comparison, and sufficiency of samples are computed (see Sec. 4.1).
@@
P-1 [83] 2012 PD Senosr+Vision
@@
P-2 [72] 2013 PD Sensor EMG SVM 10 90.0 0.33
P-3 [65] 2013 PD Sensor IMUs AdaBoost 173 81.0 0.56
P-4 [155] 2013 PD Senosr+Vision
@@
P-5 [156] 2013 PD Sensor vGRF SVM 62 81.5-83.4 0.56
P-6 [157] 2013 PD Sensor IMUs RF, DT, NB 16 96.1 0.33
P-7 [108] 2014 PD Sensor Force sensitive resistor LS-SVM, HMM 31 90.3 0.67
P-8 [64] 2014 PD Sensor IMUs RF 20 98.0 0.33
P-9 [49] 2015 PD Sensor vGRF KFD, BA, KNN, SVM, RF 49 82.0-92.6 0.44
P-10 [80] 2015 PD Vision RGB Camera NN 51 97.2 0.67
P-11 [158] 2015 PD Sensor vGRF KNN 166 83.0 0.56
P-12 [159] 2015 PD Senosr+Vision Kinect (Camera+Depth sensor)
@@
P-13 [60] 2016 PD Sensor IMUs+Force sensitive resistor LDA, SVM, ANN 166 83.3-90.0 0.56
P-14 [109] 2016 PD Sensor vGRF
@@
P-15 [142] 2016 PD Sensor Force sensitive resistor Neural networks 165 87.9 1.00
P-16 [110] 2016 PD Sensor vGRF
@@
P-17 [143] 2016 PD Sensor Force sensitive resistor Q-BTDNN 166 90.9-92.2 1.00
P-18 [160] 2017 PD Sensor IMUs KNN 580 80.8-90.2 0.56
P-19 [161] 2017 PD Sensor IMUs (Smartphone) RF 50 87.0 0.44
P-20 [57] 2017 PD Sensor IMUs+vGRF SVM, KNN, RF, DT 47 93.6 0.67
P-21 [111] 2017 PD Sensor vGRF SVM 31 90.3-100 0.67
P-22 [162] 2017 PD Sensor Walkway system (GAITRite) RF, RBF-SVM 80 85.0 0.44
P-23 [144] 2017 PD Sensor vGRF Tensor decomposition 165 100.0 0.56
P-24 [163] 2017 PD Sensor vGRF GLRA, SVM 58 82.8-84.5 0.44
P-25 [164] 2018 PD Sensor IMUs MLP, DBNs 45 93.5-94.5 0.56
P-26 [165] 2018 PD Sensor Kinect LR, NB, RF, DT 60 61.0-82.0 0.44
P-27 [50] 2018 PD Sensor vGRF SVM 166 94.8 0.56
P-28 [166] 2018 PD Sensor IMUs NB, LDA, k-NN, DT, SVM 50 68.9-75.6 0.44
P-29 [67] 2018 PD Sensor IMUs+Force sensitive resistor
@@
P-30 [112] 2018 PD Sensor IMUs SVM, RF, NB 90 78.0-96.7 0.67
P-31 [144] 2018 PD Sensor vGRF Tensor model 165 86.2 0.78
P-32 [113] 2018 PD Sensor IMUs SVM, KNN, DT, NB 51 89.1 0.67
P-33 [167] 2018 PD Sensor vGRF LR, DT, RF, SVM, KNN 166 93.1 0.56
Table 3: Continued table of Table 2 for AI-assisted PD diagnosis based on gait.
P-34 [168] 2018 PD Sensor vGRF SVM, PSO 135 87.1-95.7 0.56
P-35 [75] 2018 PD Vision Camera+Reflective markers
@@
P-36 [169] 2018 PD Sensor Kinect Fisher’s Score (FS) 40 75.0-85.0 0.33
P-37 [125] 2018 PD Sensor vGRF
@@
P-38 [114] 2018 PD Sensor vGRF RBF-SVM 64 90.6 0.67
P-39 [126] 2018 PD Sensor vGRF LSTM, CNN 166 93.9-98.6 0.89
P-40 [79] 2019 PD Vision Camera (Kinect) SVM, ANN 30 89.4 0.44
P-41 [170] 2019 PD Sensor IMUs SVM, RF 196 0.805-0.878 (AUC) 0.56
P-42 [115] 2019 PD Sensor vGRF
@@
P-43 [81] 2019 PD Vision Camera (Kinect) Conv LSTM 182 64.0-83.0 0.67
P-44 [171] 2019 PD Sensor IMUs
@@
P-45 [69] 2019 PD Sensor EEG RF, SVM, KNN 41 85.7 0.44
P-46 [172] 2019 PD Sensor IMUs CNN 84 97.6 0.56
P-47 [116] 2019 PD Sensor Force sensitive resistor SVM 166 88.9-100.0 0.78
P-48 [173] 2019 PD Sensor Walkway system (GAITRite) RF, SVM, LR 303 97.0 0.56
P-49 [174] 2019 PD Sensor+Vision
@@
P-50 [117] 2019 PD Sensor vGRF NN 166 98.8 0.78
P-51 [175] 2020 PD Sensor Force sensitive resistor
@@
P-52 [127] 2020 PD Sensor vGRF DNN 165 98.7 0.89
P-53 [66] 2020 PD Sensor IMUs+Force sensitive resistor SVM 10 97.4-98.8 0.33
P-54 [176] 2020 PD Sensor vGRF Discriminant analysis 47 95.0 0.44
P-55 [177] 2020 PD Sensor IMUs (Smart Phone) RF 456 85.5-95.0 0.56
P-56 [178] 2020 PD Sensor vGRF NN 166 97.4 0.89
P-57 [118] 2020 PD Sensor Force sensitive resistor
@@
P-58 [128] 2020 PD Sensor vGRF CNN, LSTM 166 98.0 0.89
P-59 [66] 2020 PD Sensor IMUs+Force sensitive resistor
@@
P-60 [119] 2020 PD Sensor IMUs MLP, SVM 34 68.9 0.67
P-61 [179] 2020 PD Sensor IMUs PLS-DA 93 97.9 0.44
P-62 [180] 2020 PD Sensor IMUs LR 16 93.6 0.33
P-63 [129] 2020 PD Sensor vGRF NN, SVM, KNN, DT, RF 567 49.0-61.0 (F1 score) 0.67
P-64 [130] 2020 PD Vision Camera Mask R-CNN 301 97.3 0.67
P-65 [120] 2020 PD Sensor vGRF SVM, DT 166 69.9-99.4 0.78
Table 4: Continued table of 3 for AI-assisted PD diagnosis based on gait.
P-66 [131] 2020 PD Sensor IMUs Bidirectional-LSTM 114 82.4 0.67
P-67 [132] 2020 PD Sensor IMUs CNN 2804 0.86 (AUC) 0.67
P-68 [133] 2020 PD Sensor vGRF ANN 166 98.3 0.89
P-69 [134] 2021 PD Sensor vGRF LSTM 64 98.6 0.89
P-70 [135] 2021 PD Sensor vGRF Bi-LSTM 116 91.2-100.0 0.89
P-71 [121] 2021 PD Sensor vGRF KNN, NB, SVM 165 88.5-98.8 0.78
P-72 [181] 2021 PD Sensor IMUs GMM 380 0.69 (AUC) 0.56
P-73 [182] 2021 PD Sensor IMUs
@@
P-74 [136] 2021 PD Sensor Force sensitive resistor CNN, XGBoost 166 97.3-98.4 0.78
P-75 [145] 2021 PD Sensor vGRF CNN 166 92.7-98.3 1.00
P-76 [84] 2021 PD Others Electronic medical record LR, RNN 28216 0.874 (AUC) 0.78
P-77 [73] 2021 PD Vision
@@
P-79 [183] 2021 PD Sensor vGRF LSTM 11 84.5 0.67
P-80 [184] 2021 PD Vision Camera CNN 22 94.9 0.44
P-81 [122] 2021 PD Sensor vGRF
@@
P-82 [146] 2021 PD Sensor vGRF NN 166 84.8 1.00
P-83 [185] 2021 PD Sensor IMUs SVM 300 96.7 0.56
P-84 [137] 2021 PD Sensor vGRF CNN 166 97.0 0.67
P-85 [147] 2022 PD Sensor vGRF KNN, Fuzzy neural network 165 97.6 1.00
P-86 [63] 2022 PD Sensor IMUs CNN, GRU 134 83.7-92.7 0.67
P-87 [123] 2022 PD Sensor vGRF RF 165 94.5-98.8 0.78
P-88 [186] 2022 PD Sensor vGRF LR, SVM, DT, KNN 306 81.0-85.0 0.56
P-89 [187] 2022 PD Sensor IMUs SVM 115 92.7-96.3 0.56
P-90 [188] 2022 PD Sensor IMUs SVM, RF 99 64.1-72.7 0.44
P-91 [189] 2022 PD Sensor IMUs SVM, DT, RF, KNN, MLP 128 74.0-86.0 0.56
P-92 [82] 2022 PD Vision Camera (Kinect) LR, SVM, DT, NB, RF 60 81.8-84.5 0.44
P-93 [148] 2022 PD Sensor IMUs PD-ResNet 457 95.5 1.00
P-94 [190] 2022 PD Sensor Force sensitive resistor SVM, RF 22 99.6 0.33
P-95 [191] 2022 PD Sensor IMUs (Smart phone) RF 1397 73.0 0.56
P-96 [192] 2022 PD Vision Camera
@@
P-97 [149] 2022 PD Sensor vGRF CNN, RF 166 99.5 1.00
P-98 [150] 2022 PD Vision Camera
@@
P-99 [193] 2022 PD Sensor IMUs CNN, NN 17 91.0-92.0 0.56
Table 5: Continued table of 4 for AI-assisted PD diagnosis based on gait.
P-100 [194] 2022 PD Sensor Retro-reflective markers
@@
P-101 [195] 2022 PD Sensor IMUs XGBoost, SVM, RF 134 81.0-88.3 0.56
P-102 [78] 2022 PD Sensor+Vision Camera+Force sensitive resistor AdaBoost, Extra Trees, RF 44 88.0 0.44
P-103 [76] 2022 PD Sensor+Vision Camera+Force sensitive resistor
@@
P-104 [138] 2022 PD Sensor IMUs DCNN, LSTM 30 83.3 0.78
P-105 [196] 2022 PD Sensor IMUs (Smart phone)
@@
P-106 [77] 2022 PD Sensor+Vision Camera (Kinect)+vGRF+IMUs
@@
P-107 [197] 2022 PD Vision Camera CNN, SVM 22 96.6 0.44
P-108 [198] 2022 PD Vision Camera
@@
P-109 [151] 2022 PD Vision Camera (Smart phone) SlowFast GCN network 68 77.1-87.5 0.89
P-110 [68] 2022 PD Sensor Force sensitive resistor+IMUs RF, LR, Gradient Boosting 22 97.3 0.33
P-111 [124] 2022 PD Sensor vGRF SVM 166 98.2 0.78
P-112 [199] 2022 PD Sensor vGRF SVM, ANN 48318 93.0 0.56
P-113 [46] 2022 PD Sensor+Vision
@@
P-114 [200] 2022 PD Sensor IMUs MLP, KNN, RF, SVM 30 100.0 0.56
P-115 [139] 2022 PD Sensor vGRF ANN 166 64.4-86.8 0.67
P-116 [140] 2023 PD Sensor vGRF 2D CNN 166 98.0 0.89
P-117 [201] 2023 PD Sensor IMUs
@@
P-118 [202] 2023 PD Sensor IMUs CNN 90 82.8-86.4 0.55
P-119 [203] 2023 PD Sensor IMUs PCA, SVM 280 88.0-90.0 0.55
P-120 [141] 2023 PD Sensor Force sensitive resistor
@@
P-121 [152] 2023 PD Vision Camera (Smart phone)
@@
P-122 [204] 2023 PD Sensor IMUs SVM 78 82.0 0.44
P-123 [154] 2023 PD Sensor IMUs
@@
P-124 [205] 2023 PD Sensor IMUs Non-linear model 53 84.9 0.44
Table 6: Existing studies (C-1 to C-19) that include combinatorial diagnosis of different NDs (PD, ALS, HD) based on gait. We summarize input data, main AI model(s), experimental sample size, accuracy (range), and quality score of each study. Quality scores that evaluate the novelty of AI models, comprehensiveness of method comparison, and sufficiency of experimental samples are computed (see Sec. 4.1).
@@
C-1 [101] 2012 PD, ALS, HD Sensor Force sensitive resistor SVM 64 90.6 0.44
C-2 [206] 2013 PD, ALS, HD Sensor vGRF Elman’s recurrent neural network (ERNN) 63 90.6 0.78
C-3 [102] 2015 PD, ALS, HD Sensor Force sensitive resistor SVM, RF, MLP, KNN 64 96.6-100.0 0.44
C-4 [207] 2017 PD, ALS, HD Sensor vGRF SVM, EC, DT, KNN 62 79.3 0.56
C-5 [208] 2017 PD, ALS, HD Sensor vGRF Adaboost 64 87.3-92.3 0.67
C-6 [209] 2018 PD, ALS, HD Sensor vGRF RF, SVM 64 90.6-96.9 0.44
C-7 [210] 2018 PD, ALS, HD Sensor Force sensitive resistor fusszy KNN 64 98.4 0.44
C-8 [211] 2020 PD, ALS, HD Sensor Force sensitive resistor Random forest 64 79.3-91.7 0.67
C-9 [212] 2020 PD, ALS, HD Vision Camera Discriminant Analysis, RF, Multi-SVM, KNN 64 92.2 0.44
C-10 [213] 2020 PD, ALS, HD Sensor Force sensitive resistor SVM, KNN 64 99.8-100.0 0.44
C-11 [214] 2020 PD, ALS, HD Sensor vGRF CNN 64 98.4-100.0 0.56
C-12 [103] 2020 PD, ALS, HD Sensor vGRF
@@
C-13 [215] 2020 PD, ALS, HD Sensor Force sensitive resistor KNN 64 90.5 0.44
C-14 [216] 2021 PD, ALS, HD Sensor vGRF CNN, SVM 64 97.4-100.0 0.78
C-15 [77] 2022 PD, ALS, HD Sensor+Vision
@@
C-16 [217] 2022 PD, ALS, HD Sensor vGRF AlextNet CNN 63 96.4 0.78
C-17 [218] 2022 PD, ALS, HD Sensor vGRF RF, classification trees 64 69.6-95.1 0.67
C-18 [219] 2023 PD, ALS, HD Sensor Force sensitive resistor